OpenStudy (pokiedokie):

Without graphing, classify each system of equations as independent, dependent, or inconsistent. (please help me with these two!! I'm slightly confused) {6y + 2x = 8 {12y + 4x = 4 and {2 - 0.25x = 0.5y {-1.5y = 1.5x - 3

OpenStudy (anonymous):

The easiest way to solve this is by solving the system of equations either by substitution or elimination. Let's take a look at the first one: 6y + 2x = 8 12y + 4x = 4 We can solve this by elimination if we multiply the top equation by -2 and then adding the two equations together. This will give us 0 + 0 = -12. This is not a true statement. The system if equations is inconsistent and therefore does not yield a solution. If there is no solution, then that means that these two lines are in fact parallel. Can you try to solve the second one yourself using the same process?

OpenStudy (pokiedokie):

yes, but do I have to find a pattern for that one as well like with the first equation (being it multiplied by -2)?

OpenStudy (anonymous):

In a manner of speaking yes. When solving by elimination, you want to multiply by a factor so that if you add the two equations together, either the x's or the y's cancel. In the first set, multiplying by -2 and adding cancelled out both the x's AND the y's.

OpenStudy (pokiedokie):

im sorry for being a pain, but i would rearrange this equation so I can subtract/eliminate it, correct?

OpenStudy (anonymous):

Yes, it would certainly make it easier :) 0.5y = 2- 0.25x -1.5y = 1.5x -3 It seems easier to cancel out the y. You don't have to multiply by a negative because they are already opposite signs. Try to do the rest and let me know if you're still having trouble :)

Directrix (directrix):

Study tool for later. @pokiedokie

OpenStudy (pokiedokie):

thank you very much!! this is helpful

OpenStudy (pokiedokie):

@CShrix i got -1y = -1.25-1, did i mess up somewhere because this doesn't seem correct to me?

OpenStudy (anonymous):

You have to multiply the first equation by 3 first. Then when you add the two equations, there are no more y's in the equation. Then solve for x from there :)

OpenStudy (pokiedokie):

1.5y = 6 - 0.75x -1.5y - 3 - 1.5x ________________ y = 3 - (-0.75) ??

OpenStudy (anonymous):

1.5y-1.5y = 0. So you have 0 = 3 + 0.75x

OpenStudy (pokiedokie):

whoops, sorry, so this would be inconsistent

OpenStudy (pokiedokie):

thank you so much!!